solve the system of equations by letting A 1/X and B = 1/y. (If there is no solution, enter NO SOLUTION. If the system has an infinite number of solutions, set y t and solve torx in terms of t.) 20 20 (x, y) =
Answer #1

20A-20B=9 (1)

4A+5B=0. (2)

Multiplying equation 2 by 5

20A+25B=0. (3)

(3) - (1) gives. 45B= -9

B=rac{-9}{45}=rac{-1}{5}

Putting value of B in (1) gives

20A=9+20B=9-4=5

A=- 20 4

A=1/x gives X=4

B=1/y gives Y= -5